The Millennium Square Inn’s total room revenue is generated by its suite and standard rooms. Total revenue generated by the suit room sales is $95,000 in January. Also, total revenue generated by the standard room sales is $67,000 in the same month. The management forecasted that there will be 8.00% increase in revenues for suites and 6.00% decrease in revenues for standard rooms in February. Based on the information given, calculate the projected total room revenue for the Millennium Square Inn in February.
Answer #1

Calculate total revenue

Total revenue Standard room Suite
January 95000 67000 95000-67000 = 28000
February 62980+30240 = 93220 67000*94% = 62980 28000*1.08 = 30240

So total room revenue for Millennium Square in February = $93220